# Skill-aware Mutual Information Optimisation for Generalisation in Reinforcement Learning

### v1

@@ Abstract
Meta-Reinforcement Learning (Meta-RL) agents can struggle to operate across tasks with varying environmental features that require different optimal skills (i.e., different modes of behaviours). Using context encoders based on contrastive learning to enhance the generalisability of Meta-RL agents is now widely studied but faces challenges such as the requirement for a large sample size, also referred to as the $\log$-$K$ curse. To improve RL generalisation to different tasks, we first introduce Skill-aware Mutual Information (SaMI), an optimisation objective that aids in distinguishing context embeddings according to skills, thereby equipping RL agents with the ability to identify and execute different skills across tasks. We then propose Skill-aware Noise Contrastive Estimation (SaNCE), a $K$-sample estimator used to optimise the SaMI objective. We provide a framework for equipping an RL agent with SaNCE in practice and conduct experimental validation on modified MuJoCo and Panda-gym benchmarks. We empirically find that RL agents that learn by maximising SaMI achieve substantially improved zero-shot generalisation to unseen tasks. Additionally, the context encoder equipped with SaNCE demonstrates greater robustness to reductions in the number of available samples, thus possessing the potential to overcome the $\log$-$K$ curse.
@@ 1 Introduction
Meta-Reinforcement Learning (Meta-RL) agents often learn policies that do not generalise across tasks in which the environmental features and optimal skills are different [des Combes et al., 2018, Garcin et al., 2024]. Consider, for example, a set of cube-moving tasks, in which an agent is required to move a cube to a goal position on a table (Figure 1). These tasks become challenging if the underlying environmental features vary between different tasks, such as table friction. Suppose an agent
@@
attempts to push a cube across a table covered by a tablecloth and receives environmental feedback showing that the cube has become "unpushable". In this scenario, a generalist agent should infer that the table friction is relatively high. Consequently, the agent ought to adapt its strategy, opting to lift the cube off the table to avoid friction, rather than continuing to push. Recent advances in Meta-RL [Lee et al., 2020, Agarwal et al., 2021, Mu et al., 2022, Dunion et al., 2023b, a] have shown promising potential to perceive and understand environmental features by inferring context embeddings from a small number of interactions in the environment. The context embedding is expected to capture the distribution of tasks and efficiently infer new tasks. Meta-RL methods then train a policy conditioned on the context embedding to generalise to multiple tasks.
@@
The context encoder is the key component for capturing the context embedding from recent experiences [Clavera et al., 2019a, Lee et al., 2020], which affects generalisation performance significantly. Some Meta-RL algorithms [Fu et al., 2021, Wang et al., 2021, Li et al., 2021, Sang et al., 2022] are equipped with context encoders based on contrastive learning, which uses the InfoNCE lower bound [Oord et al., 2019] to optimise the mutual information (MI) between context embeddings. InfoNCE is a $K$-sample MI estimator and is utilised to learn distinct context embeddings for each task. Despite significant advancements, integrating contrastive learning with Meta-RL poses several unresolved challenges, of which two are particularly relevant to this research: (i) existing context encoders based on contrastive learning do not distinguish tasks that require different skills; many prior algorithms only pull embeddings of the same tasks together and push those of different tasks apart. However, for example, a series of cube-moving tasks with high friction may only require a Pick&Place skill (picking the cube off the table and placing it at the goal position), making further differentiation unnecessary. (ii) K-sample MI estimators are sensitive to the sample size K (i.e., the log-K curse) [Poole et al., 2019]; a substantial quantity of negative samples is required to approximate the true MI, which is challenging due to RL’s low sample efficiency [Franke et al., 2021] and often impractical for achieving accurate MI estimation [Arora et al., 2019, Nozawa and Sato, 2021]. The effectiveness of $K$-sample MI estimators breaks down with a small sample size and leads to a significant performance drop in downstream RL tasks [Mnih and Teh, 2012, Guo et al., 2022].
@@
To enhance RL generalisation across different tasks, we propose that the context embeddings should optimise downstream tasks and indicate whether the current skill remains optimal or requires further exploration, thereby addressing issue (i). This approach also reduces the necessary sample size and helps to overcome issue (ii) by concentrating solely on extracting task-relevant MI. Specifically, we propose a three-step process tailored to RL: (1) We introduce Skill-aware Mutual Information (SaMI), a smaller ground-truth MI that discriminates context embeddings according to skills by maximising the MI between context embedding, skills and trajectories. Additionally, we provide a theoretical proof of why introducing skills can make the ground-truth MI smaller, thus making it easier to optimise; (2) We propose a more data-efficient $K$-sample estimator, Skill-aware Noise Contrastive Estimation (SaNCE), used to optimise SaMI and reduce the negative sample space based on skills to overcome the $log$-$K$ curse; (3) We provide a framework to equip any Meta-RL algorithm with SaNCE in practice, and propose a practical skill-aware sampling method for SaNCE when the skills are not known in advance.
@@
We demonstrate empirically in MuJoCo [Todorov et al., 2012] and Panda-gym [Gallouédec et al., 2021] that SaMI improves the zero-shot generalisation performance in sets of previously unseen tasks with moderate and extreme difficulty. In the MuJoCo benchmark, SaMI enhances two Meta-RL algorithms [Yu et al., 2020, Fu et al., 2021] by achieving an average increase of 42.27% in returns for moderate testing tasks and a 53.14% improvement for extreme testing tasks. Meanwhile, in the Panda-gym benchmark, SaNCE boosts success rates by an average of 20.23% in moderate tests and 18.36% in extreme tests. This indicates SaMI’s advantage in encoding information that enables an agent to execute effective skills in downstream control tasks. SaNCE-based RL algorithms utilise smaller sample spaces while achieving improved downstream control performance, indicating their potential to overcome the $\log$-$K$ curse. Visualisation of the learned context embeddings shows distinct clusters corresponding to different skills, indicating that SaMI is a step towards more robust and versatile RL agents.
Figure 2: A policy $\pi$ conditioned on a fixed context embedding $c$ is defined as a skill $\pi(\cdot|c)$. Executing different skills (e.g., Push skill $\pi(\cdot|c_{1})$ and Pick&Place skill $\pi(\cdot|c_{2})$) shows different modes of behaviours in trajectories (e.g., $\tau^{+}_{c_{1}}$ where the cube moves on the table; $\tau^{+}_{c_{2}}$ where the cube leaves the table).
@@ 2 Related works
Meta-RL. Meta-RL methods train an agent conditioned on context embeddings to improve generalisation to unseen tasks. As the key component of Meta-RL, the quality of the context embedding can significantly affect the agent’s performance. Existing algorithms can be categorised into three types based on different context embeddings. In the first category, the context embedding is learned by minimising the downstream RL loss [Rakelly et al., 2019, Yu et al., 2020]. PEARL [Rakelly et al., 2019] learns probabilistic context embeddings by recovering the value function. Multi-task SAC + TE (TESAC) [Yu et al., 2020] uses the Task Embedding (TE) [Hausman et al., 2018] as input to policies, parameterising the learned policies via a shared embedding space and maximising average returns. However, the update signals from the RL loss are stochastic and weak, and may not capture the similarity relations among tasks [Fu et al., 2021]. The second category involves learning context embeddings through dynamics prediction [Lee et al., 2020, Zhou et al., 2019], which can make the context embeddings noisy, as they may model irrelevant dependencies and overlook task-specific information [Fu et al., 2021]. The third category employs contrastive learning [Fu et al., 2021, Wang et al., 2021, Li et al., 2021, Sang et al., 2022], achieving significant improvements in context learning. However, these methods overlook the similarity of skills between different tasks, thus failing to achieve effective zero-shot generalisation by executing different skills. Our improvements build upon this third category by distinguishing context embeddings according to different optimal skills.
Contrastive learning. Contrastive learning has been applied to RL due to its significant momentum in representation learning in recent years, attributed to its superior effectiveness [Tishby and Zaslavsky, 2015, Hjelm et al., 2019, Dunion and Albrecht, 2024], ease of implementation [Oord et al., 2019], and strong theoretical connection to mutual information (MI) estimation [Poole et al., 2019]. MI is often estimated through noise-contrastive estimation (NCE) [Gutmann and Hyvärinen, 2010], InfoNCE [Oord et al., 2019], and variational objectives [Hjelm et al., 2019]. InfoNCE has garnered recent interest due to its lower variance [Song and Ermon, 2020] and superior performance in downstream tasks. However, InfoNCE may underestimate true MI because it is constrained by the number of samples $K$. To address this issue, CCM [Fu et al., 2021] leverages InfoNCE to learn the context embedding with a large number of samples. DOMINO [Mu et al., 2022] reduces the total MI by introducing an independence assumption; however, this results in biased information. Correspondingly, we focus on proposing an unbiased alternative MI objective and a more data-efficient $K$-sample estimator tailored for downstream RL tasks, which, to our knowledge, have not been addressed in previous research.
@@ 3 Preliminaries
Reinforcement learning. In Meta-RL, we assume an environment is a distribution $\xi(e)$ of tasks $e$ (e.g. uniform in our experiments). Each task $e\sim\xi(e)$ has a similar structure that corresponds to a Markov Decision Process (MDP) [Puterman, 2014], defined by $\mathcal{M}_{e}=(\mathcal{S},\mathcal{A},R,P_{e},\gamma)$, with a state space $\mathcal{S}$, an action space $\mathcal{A}$, a reward function $R(s_{t},a_{t})$ where $s_{t}\in\mathcal{S}$ and $a_{t}\in\mathcal{A}$, state transition dynamics $P_{e}(s_{t+1}|s_{t},a_{t})$, and a discount factor $\gamma\in[0,1)$. In order to address the problem of zero-shot generalisation, we consider the transition dynamics $P_{e}(s_{t+1}|s_{t},a_{t})$ vary across tasks $e\sim\xi(e)$ according to multiple environmental features $e=\{e^{0},e^{1},...,e^{N}\}$ that are not included in states $s$ and can be continuous random variables, such as mass and friction, or discrete random variables, such as the cube’s material. For instance, in a cube-moving environment (Figure 1), an agent has different tasks that are defined by different environmental features (e.g., mass and friction). The Meta-RL agent’s goal is to learn a generalisable policy $\pi$ that is robust to such dynamic changes. Specifically, given a set of training tasks $e$ sampled from $\xi_{\text{train}}(e)$, we aim to learn a policy that can maximise the discounted returns, $\arg\max_{\pi}\mathbb{E}_{e\sim\xi_{\text{train}}(e)}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})|a_{t}\sim\pi(a_{t}|s_{t}),s_{t+1}\sim P_{e}(s_{t+1}|s_{t},a_{t})]$, and can produce accurate control for unseen test tasks $e$ sampled from $\xi_{\text{test}}(e)$.
Contrastive learning. A Meta-RL agent first consumes a collected trajectory $\tau_{t:t+T}=\{s_{t},a_{t},r_{t},s_{t+1},...,s_{t+T},a_{t+T},r_{t+T}\}$ and outputs a context embedding $c$ through a context encoder $\psi(c|\tau_{t:t+T})$, then executes policy $\pi$ conditioned on the current state $s_{t}$ and context embedding $c$ to work across multiple tasks. As a key component, a context encoder $\psi$ captures the unknown context $c$ of the task. We focus on a context encoder that optimises the InfoNCE bound $I_{\text{InfoNCE}}(x;y)$, which is a $K$-sample estimator of the MI $I(x;y)$ [Oord et al., 2019]. Given a query $x$ and a set $Y=\{y_{1},...,y_{K}\}$ of $K$ random samples containing one positive sample $y_{1}$ and $K-1$ negative samples from the distribution $p(y)$, $I_{\text{InfoNCE}}(x;y)$ is obtained by comparing pairs sampled from the joint distribution $x,y_{1}\sim p(x,y)$ to pairs $x,y_{k}$ built using a set of negative examples $y_{2:K}$:
@@
InfoNCE constructs a formal lower bound to the MI, i.e., $I_{\text{InfoNCE}}(x;y|\psi,K)\leq I(x;y)$ [Guo et al., 2022, Chen et al., 2021]. Given two inputs $x$ and $y$, their embedding similarity is $f_{\psi}(x,y)=e^{\psi(x)^{\top}\cdot\psi(y_{1})/\beta}$, where $\psi$ is the context encoder that projects $x$ and $y$ into the context embedding space, the dot product is used to calculate the similarity score between $\psi(x),\psi(y)$ pairs [Wu et al., 2018, He et al., 2020], and $\beta$ is a temperature hyperparameter that controls the sensitivity of the product. Some previous Meta-RL methods [Lee et al., 2020, Mu et al., 2022] learn a context embedding $c$ by maximising $I_{\text{InfoNCE}}(c;\tau_{c}|\psi,K)$ between the context $c$ embedded from a trajectory in the current task, and the historical trajectories $\tau_{c}$ under the same environmental features setting.
@@ 4 Skill-aware mutual information optimisation for Meta-RL
4.1 The $log$-$K$ curse of $K$-sample MI estimators
In this section, we provide a theoretical analysis of the challenge inherent in learning a $K$-sample estimator for MI, commonly referred to as the $log$-$K$ curse. Based on this theoretical analysis, we give insights to overcome this challenge.
If we have sufficient training epochs for the context encoder $\psi$ in Equation 1 and a sample size $K$, $I_{\text{InfoNCE}}(x;y|\psi,K)\approx\log K$ [Mnih and Teh, 2012, Guo et al., 2022]. Hence, the MI we can optimise is bottlenecked by the number of available samples, formally expressed as:

### v3

@@ Abstract
Meta-Reinforcement Learning (Meta-RL) agents can struggle to operate across tasks with varying environmental features that require different optimal skills (i.e., different modes of behaviour). Using context encoders based on contrastive learning to enhance the generalisability of Meta-RL agents is now widely studied but faces challenges such as the requirement for a large sample size, also referred to as the $\log$-$K$ curse. To improve RL generalisation to different tasks, we first introduce Skill-aware Mutual Information (SaMI), an optimisation objective that aids in distinguishing context embeddings according to skills, thereby equipping RL agents with the ability to identify and execute different skills across tasks. We then propose Skill-aware Noise Contrastive Estimation (SaNCE), a $K$-sample estimator used to optimise the SaMI objective. We provide a framework for equipping an RL agent with SaNCE in practice and conduct experimental validation on modified MuJoCo and Panda-gym benchmarks. We empirically find that RL agents that learn by maximising SaMI achieve substantially improved zero-shot generalisation to unseen tasks. Additionally, the context encoder trained with SaNCE demonstrates greater robustness to a reduction in the number of available samples, thus possessing the potential to overcome the $\log$-$K$ curse.
@@ 1 Introduction
Reinforcement Learning (RL) agents often learn policies that do not generalise across tasks in which the environmental features and optimal skills are different [des Combes et al., 2018, Garcin et al., 2024]. Consider a set of cube-moving tasks where an agent is required to move a cube to a goal position on a table (Figure 1). These tasks become challenging if environmental features, such as table friction, vary between tasks. When facing an unknown environment, the agent
@@
needs to explore effectively, understand the environment, and adjust its behaviour accordingly within an episode. For instance, if the agent tries to push a cube across a table covered by a tablecloth and finds it “unpushable,” it should infer that the table friction is relatively high and adapt by lifting the cube to avoid friction, rather than continuing to push. Recent advances in Meta-Reinforcement Learning (Meta-RL) [Lee et al., 2020, Agarwal et al., 2021, Mu et al., 2022, Dunion et al., 2023b, a, McInroe et al., 2024] enable agents to understand environmental features by inferring context embeddings from a small amount of exploration, and to train a policy conditioned on the context embedding to generalise to novel tasks.
@@
In recent years, unsupervised contrastive learning algorithms have been shown to learn context embeddings that perform remarkably well on generalisation tasks [Clavera et al., 2019a, Lee et al., 2020]. In particular, some Meta-RL algorithms [Fu et al., 2021, Wang et al., 2021, Li et al., 2021, Sang et al., 2022] train context encoders by maximising InfoNCE [Oord et al., 2019], which has yielded vital insights into contrastive learning through the lens of mutual information (MI) analysis. The InfoNCE objective can be interpreted as a $K$-sample lower bound on the MI between trajectories and context embeddings. Despite significant advances, integrating contrastive learning with Meta-RL poses several unresolved challenges, of which two are particularly relevant to this research: (i) Existing context encoders based on contrastive learning do not distinguish tasks that require different skills. Many prior algorithms only pull embeddings of the same tasks together and push those of different tasks apart. However, for example, a series of cube-moving tasks with high friction may only require a Pick&Place skill (picking the cube off the table and placing it at the goal position), making further differentiation unnecessary. (ii) Existing K-sample MI estimators are sensitive to the sample size K (i.e., the log-K curse) [Poole et al., 2019]. The low sample efficiency of RL [Franke et al., 2021] and the sample limitations in zero-shot generalisation make collecting a substantial quantity of samples often impractical [Arora et al., 2019, Nozawa and Sato, 2021]. The effectiveness of $K$-sample MI estimators breaks down with a finite sample size and leads to a significant performance drop in downstream RL tasks [Mnih and Teh, 2012, Guo et al., 2022].
@@
To enhance RL generalisation across different tasks, we propose that the context embeddings should optimise downstream tasks and indicate whether the current skill remains optimal or requires further exploration. This also reduces the necessary sample size by focusing solely on skill-related information. In this work, (1) we introduce Skill-aware Mutual Information (SaMI), a generalised form of MI objective between context embeddings, skills, and trajectories, designed to address issue (i). We provide a theoretical proof showing that by introducing skills as a third variable into the MI of context embeddings and trajectories, the resulting SaMI is smaller and easier to optimise. Furthermore, (2) we propose a data-efficient $K$-sample estimator, Skill-aware Noise Contrastive Estimation (SaNCE) to optimise SaMI, effectively addressing issue (ii). Additionally, (3) we propose a practical skill-aware trajectory sampling strategy that shows how to sample positive and negative examples without relying on any prior skill distribution. In that way, Meta-RL agents autonomously acquire a set of skills applicable to many tasks, with these skills emerging solely from the SaMI learning objective and data.
@@
We demonstrate empirically in MuJoCo [Todorov et al., 2012] and Panda-gym [Gallouédec et al., 2021] that SaMI enhances the zero-shot generalisation capabilities of two Meta-RL algorithms [Yu et al., 2020, Fu et al., 2021] by achieving higher returns and success rates on previously unseen tasks, ranging from moderate to extreme difficulty. Visualisation of the learned context embeddings reveals distinct clusters corresponding to different skills, suggesting that the SaMI learning objective enables the context encoder to capture skill-related information from trajectories and incentivise Meta-RL agents to acquire a diverse set of skills. Moreover, SaNCE enables Meta-RL algorithms to use smaller sample spaces while achieving improved downstream control performance, indicating their potential to overcome the $\log$-$K$ curse.
Figure 2: A policy $\pi$ conditioned on a fixed context embedding $c$ is defined as a skill $\pi(\cdot|c)$ (shortened as $\pi_{c}$). The policy $\pi$ conditioned on a fixed $c$ alters the state of the environment in a consistent way, thereby exhibiting a mode of skill. The skill $\pi(\cdot|c_{1})$ moves the cube on the table in trajectory $\tau^{+}_{c_{1}}$ and is referred to as the Push skill; correspondingly, the Pick&Place skill $\pi(\cdot|c_{2})$ takes the cube off the table and places it in the goal position in the trajectory $\tau^{+}_{c_{2}}$.
@@ 2 Related works
Meta-RL. By conditioning on an effective context embedding, Meta-RL policies can zero-shot generalise to new tasks with a small amount of exploration Kirk et al. [2023]. Existing algorithms can be categorised into three types based on different context embeddings. In the first category, the context embedding is learned by minimising the downstream RL loss [Rakelly et al., 2019, Yu et al., 2020]. PEARL [Rakelly et al., 2019] learns probabilistic context embeddings by recovering the value function. Multi-task SAC with task embeddings as inputs to policies (TESAC) [Hausman et al., 2018, Yu et al., 2020] parameterises the learned policies through a shared embedding space, aiming to maximise the average returns across tasks. However, the update signals from the RL loss are stochastic and weak, and may not capture the similarity relations among tasks [Fu et al., 2021]. The second category involves learning context embeddings through dynamics prediction [Lee et al., 2020, Zhou et al., 2019], which can make the context embeddings noisy, as they may model irrelevant dependencies and overlook task-specific information [Fu et al., 2021]. The third category employs contrastive learning [Fu et al., 2021, Wang et al., 2021, Li et al., 2021, Sang et al., 2022], achieving significant improvements in context learning. However, these methods overlook the similarity of skills between different tasks, thus failing to achieve effective zero-shot generalisation by executing different skills. Our improvements build upon this third category by distinguishing context embeddings according to different optimal skills.
Contrastive learning. Contrastive learning has been applied to RL due to its significant momentum in representation learning in recent years, attributed to its superior effectiveness [Tishby and Zaslavsky, 2015, Hjelm et al., 2019, Dunion and Albrecht, 2024], ease of implementation [Oord et al., 2019], and strong theoretical connection to MI estimation [Poole et al., 2019]. MI is often estimated using InfoNCE [Oord et al., 2019] that has gained recent attention due to its lower variance [Song and Ermon, 2020] and superior performance in downstream tasks. However, InfoNCE may underestimate the true MI when the sample size $K$ is finite. To address this limitation, CCM [Fu et al., 2021] uses a large number of samples for maximising InfoNCE. DOMINO [Mu et al., 2022] reduces the true MI by introducing an independence assumption; however, this results in biased estimates. We focus on proposing an unbiased alternative MI objective and a more data-efficient $K$-sample estimator tailored for downstream RL tasks, which, to our knowledge, have not been addressed in previous research.
@@ 3 Preliminaries
Reinforcement learning. In Meta-RL, we assume an environment is a distribution $\xi(e)$ of tasks $e$ (e.g. uniform in our experiments). Each task $e\sim\xi(e)$ has a similar structure that corresponds to a Markov Decision Process (MDP) [Puterman, 2014], defined by $\mathcal{M}_{e}=(\mathcal{S},\mathcal{A},R,P_{e},\gamma)$, with a state space $\mathcal{S}$, an action space $\mathcal{A}$, a reward function $R(s_{t},a_{t})$ where $s_{t}\in\mathcal{S}$ and $a_{t}\in\mathcal{A}$, state transition dynamics $P_{e}(s_{t+1}|s_{t},a_{t})$, and a discount factor $\gamma\in[0,1)$. In order to address the problem of zero-shot generalisation, we consider the transition dynamics $P_{e}(s_{t+1}|s_{t},a_{t})$ vary across tasks $e\sim\xi(e)$ according to multiple environmental features $e=\{e^{0},e^{1},...,e^{N}\}$ that are not included in states $s$ and can be continuous random variables, such as mass and friction, or discrete random variables, such as the cube’s material. For instance, in a cube-moving environment (Figure 1), an agent has different tasks that are defined by different environmental features (e.g., mass and friction). The Meta-RL agent’s goal is to learn a generalisable policy $\pi$ that is robust to such dynamic changes. Specifically, given a set of training tasks $e$ sampled from $\xi_{\text{train}}(e)$, we aim to learn a policy that can maximise the discounted returns, $\arg\max_{\pi}\mathbb{E}_{e\sim\xi_{\text{train}}(e)}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})|a_{t}\sim\pi(a_{t}|s_{t}),s_{t+1}\sim P_{e}(s_{t+1}|s_{t},a_{t})]$, and can produce accurate control for unseen test tasks sampled from $\xi_{\text{test}}(e)$.
Contrastive learning. In Meta-RL, the context encoder $\psi(c|\tau_{c,0:t})$ first takes the trajectory $\tau_{c,0:t}=\{s_{0},a_{0},r_{0},...,s_{t}\}$ from the current episode as input and compresses it into a context embedding $c$ [Fu et al., 2021]. Then, the policy $\pi$, conditioned on context embedding $c$, consumes the current state $s_{t}$ and outputs the action $a_{t}$. As a key component, the quality of context embedding $c$ can affect algorithms’ performance significantly. MI is an effective measure of embedding quality [Goldfeld et al., 2019], hence we focus on a context encoder that optimises the InfoNCE objective $I_{\text{InfoNCE}}(x;y)$, which is a $K$-sample estimator and lower bound of the MI $I(x;y)$ [Oord et al., 2019]. Given a query $x$ and a set $Y=\{y_{1},...,y_{K}\}$ of $K$ random samples containing one positive sample $y_{1}$ and $K-1$ negative samples from the distribution $p(y)$, $I_{\text{InfoNCE}}(x;y)$ is obtained by comparing pairs sampled from the joint distribution $x,y_{1}\sim p(x,y)$ to pairs $x,y_{k}$ built using a set of negative examples $y_{2:K}$:
@@
InfoNCE constructs a formal lower bound on the MI, i.e., $I_{\text{InfoNCE}}(x;y|\psi,K)\leq I(x;y)$ [Guo et al., 2022, Chen et al., 2021]. Given two inputs $x$ and $y$, their embedding similarity is $f_{\psi}(x,y)=e^{\psi(x)^{\top}\cdot\psi(y)/\beta}$, where $\psi$ is the context encoder that projects $x$ and $y$ into the context embedding space, the dot product is used to calculate the similarity score between $\psi(x),\psi(y)$ pairs [Wu et al., 2018, He et al., 2020], and $\beta$ is a temperature hyperparameter that controls the sensitivity of the product. Some previous Meta-RL methods [Lee et al., 2020, Mu et al., 2022] learn a context embedding $c$ by maximising $I_{\text{InfoNCE}}(c;\tau_{c}|\psi,K)$ between the context $c$ embedded from a trajectory in the current task, and the historical trajectories $\tau_{c}$ under the same environmental features setting.
@@ 4 Skill-aware mutual information optimisation for Meta-RL
4.1 The $\boldsymbol{\log}$-$\boldsymbol{K}$ curse of $\boldsymbol{K}$-sample MI estimators
In this section, we provide a theoretical analysis of the challenge inherent in learning a $K$-sample estimator for MI, commonly referred to as the $\log$-$K$ curse. Based on this theoretical analysis, we give insights to overcome this challenge. Given that we focus on the generalisation of RL, we only consider cases with a finite sample size of $K$. If a context encoder $\psi$ in Equation (1) has sufficient training epochs, then $I_{\text{InfoNCE}}(x;y|\psi,K)\approx\log K$ [Mnih and Teh, 2012, Guo et al., 2022]. Hence, the MI we can optimise is bottlenecked by the number of available samples, formally expressed as: